### v2

# Vertex-minimal hyperbolic origami 2-torus

Abstract
We show that there exists a geodesic triangulation $\mathcal{T}$ of a hyperbolic genus 2 surface $\Sigma_{2}$ with 10 vertices and an isometric polyhedral embedding $S:\Sigma_{2}\hookrightarrow\mathbb{H}^{3}$ that sends the triangles in $\mathcal{T}$ to geodesic triangles in $\mathbb{H}^{3}$. We call this type of embedding a hyperbolic origami 2-torus. Since 10 is the combinatorially minimum number of vertices required to triangulate a genus 2 surface, this paper settles the question of minimum number of vertices required to obtain a hyperbolic origami 2-torus.
1 Introduction
1.1 Context and Problem Statement
A hyperbolic 2-torus $\Sigma_{2}$ is a quotient of $\mathbb{H}^{2}$ by the action of a surface group $\Gamma<\operatorname{Isom}^{+}(\mathbb{H}^{2})$ such that $\mathbb{H}^{2}/\Gamma$ is a genus 2 surface. Here, $\Gamma$ is isomorphic to the group with presentation $\langle a,b,c,d\,|\,[a,b][c,d]\rangle$, where $a,b]=aba^{-1}b^{-1}$ denotes the commutator of the two elements, and the action of $\Gamma$ on $\mathbb{H}^{2}$ is free and properly discontinuous. Every hyperbolic 2-torus $\Sigma_{2}$ inherits the structure of a Riemannian manifold of constant sectional curvature -1 from $\mathbb{H}^{2}$. A hyperbolic origami 2-torus is a map $S:(\Sigma_{2},\mathcal{T})\hookrightarrow\mathbb{H}^{3}$, where $\mathcal{T}$ is a geodesic triangulation of $\Sigma_{2}$, and $S$ restricts to isometries on the triangles of $\mathcal{T}$. Concretely, we think of a hyperbolic origami 2-torus $S$ as gluing hyperbolic triangles together in $\mathbb{H}^{3}$, where the cone angles at each vertex are $2\pi$.
The study of graph embeddings and triangulations of closed surfaces can be dated back to the late 19th century [Hea90]. The minimum number of vertices $n_{0}$ to triangulate a closed surface of Euler characteristic $e$ is given by
$n_{0}\geq\bigg\lfloor\frac{7+\sqrt{49-24e}}{2}\bigg\rfloor$ (1)
Equation (1) is proved by Ringel [Rin55] for the non-orientable cases and later by Jungerman and Ringel [JR80] for the orientable cases. However, the bound in Equation (1) is not tight for the 2-torus. In [Hun78], Huneke showed that the minimum number of vertices required to triangulate a 2-torus is 10, in contrast with a looser lower bound of 8 from Equation (1). In 2008, Lutz [Lut08] enumerated all 865 isomorphism classes of 10-vertex triangulations of the 2-torus.
Geometers were interested in whether these triangulated surfaces admit piecewise-affine embeddings into $\mathbb{R}^{3}$ that are affine on the triangle pieces [BB87, Bre87, BB89, BZ95, HLZ06]. These embeddings are called polyhedral embeddings. A polyhedral embedding of a 10-vertex triangulation of the 2-torus was first discovered by Brehm [Bre81]. Later, Hougardy, Lutz, and Zelke [HLZ07] developed a computer algorithm to enumerate polyhedral embeddings of all 10-vertex triangulations of the 2-torus. However, a fundamental problem remains unsolved: if we take into account the hyperbolic structures on a 2-torus, these polyhedral embeddings fail to preserve them. First, it’s impossible to map hyperbolic triangles isometrically to geodesic triangles in $\mathbb{R}^{3}$ due to the Gauss–Bonnet theorem. Thus, it is more natural to consider embeddings of $(\Sigma_{2},\mathcal{T})$ into $\mathbb{H}^{3}$, where the images of triangles in $\mathcal{T}$ are isometric geodesic triangles in the hyperbolic 3-space.
There has been significant progress on polyhedral embeddings of the flat torus. See [Zal00, BH08, Seg16, LT24, Tsu24]. A flat polyhedral torus is a piecewise-linear embedding of a triangulated torus $T^{2}$ into $\mathbb{R}^{3}$ that restricts to linear isometries on the triangles, where the triangles in $T^{2}$ come from a triangular tiling of the universal cover $\mathbb{R}^{2}$. A very recent work of Schwartz [Sch25] showed that a flat polyhedral torus of 7 vertices doesn’t exist, but one can obtain one with 8 vertices with 2-fold symmetry. This settled the question of the minimum number of vertices required for a flat polyhedral torus. In this paper, we prove that the minimum number of vertices required for a hyperbolic origami 2-torus is 10.
1.2 The Candidate Model 2-Torus
We state the main result of this paper.
Theorem 1.1.
There exists a hyperbolic 2-torus $\Sigma_{2}$ along with a 10-vertex geodesic triangulation $\mathcal{T}$ and a hyperbolic origami 2-torus $S:(\Sigma_{2},\mathcal{T})\hookrightarrow\mathbb{H}^{3}$.
Figure 1: The candidate 2-torus $\hat{S}$ in the Beltrami–Klein model $\mathbb{H}^{3}$, shown from opposite sides of a “hole” passing through it. Triangles are shaded along a yellow–purple spectrum according to the $y$-coordinates of their centroids, ordered from highest to lowest.
Figure 2: Left: The induced triangle tiling from $\mathcal{T}$ in the Beltrami–Klein model of $\mathbb{H}^{2}$. Right: The tiling in the Poincaré disk model. The colors of the triangles are derived from Figure 1.
The information of a hyperbolic origami 2-torus is captured entirely by the triangulation $\mathcal{T}$ and the image of the vertices of $\mathcal{T}$ under the map $S:(\Sigma_{2},\mathcal{T})\rightarrow\mathbb{H}^{3}$, as long as the cone angles of the vertices of $\mathcal{T}$ are all $2\pi$. We remark that Theorem 1.1 implies the existence of hyperbolic origami 2-tori with $n$ vertices for all $n\geq 10$; one simply uses the 10-vertex embedding and adds vertices to the interior of the triangles in $\mathcal{T}$ to obtain triangulations with more vertices.
Consider the following 10-vertex triangulation $\mathcal{T}=(V,E,F)$ with vertices in $V$, edges in $E$, and oriented triangles in $F$. We index $V$ by $\{0,\ldots,9\}$. We list the degree sequence of the 10 vertices in $V$:
[…]
We also list out the indices of the 24 oriented triangles in $F$ in the form of 3-tuples. A tuple $(i,j,k)$ indicates an oriented face joined by vertices $i,j,k$.
$\begin{matrix}(0,1,7)&(0,2,1)&(0,3,6)&(0,4,9)&(0,5,3)&(0,6,4)\\ (0,7,8)&(0,8,5)&(0,9,2)&(1,2,4)&(1,3,7)&(1,4,6)\\ (1,5,8)&(1,6,5)&(1,8,3)&(2,3,8)&(2,6,3)&(2,7,4)\\ (2,8,7)&(2,9,6)&(3,4,7)&(3,5,4)&(4,5,9)&(5,6,9)\end{matrix}$ (2)
The proof of Theorem 1.1 consists of two steps. First, we find a polyhedral 2-torus $\hat{S}$ that is “robustly embedded” and “almost isometric.” Next, we argue that an isometric one exists by perturbing the coordinates of $\hat{S}$. Figure 1 shows the polyhedral 2-torus $\hat{S}$ embedded in the Beltrami–Klein model of the hyperbolic 3-space (recall that in the Klein model, geodesics are straight lines and geodesic triangles are Euclidean triangles). In Figure 2, we draw the induced tiling of the universal cover $\mathbb{H}^{2}$ by the “almost geodesic” triangulation on $\Sigma_{2}$ by “almost isometrically” lifting the triangles in $\hat{S}$ to $\mathbb{H}^{2}$ and gluing them together. Since $\hat{S}$ is not a hyperbolic origami 2-torus, the picture of the tiling in the universal cover is only meant to be an approximation up to some error (about $10^{-28}$), but the error is too small to be discernible by human eyes so that visually the picture looks reasonable. See Figure 3, 4, 5 for the intersection of $\hat{S}$ with the $xy$-plane, xzxz-plane, and $yz$-plane.
Figure 3: The intersection of $\hat{S}$ with the $xy$-plane, which is the boundary of a simply-connected component colored in yellow.
Here we give the exact coordinates of the vertices $\hat{X}_{0},\ldots,\hat{X}_{9}$ of our candidate 2-torus $\hat{S}$ in the Beltrami–Klein model:
$\begin{matrix}0.7315&0.0202&0.28688022781563440615364787558404\\ -0.316&0.5792&-0.2252919753182150895621576631383\\ 0.3426&-0.592&-0.22851917827575874828458055465199\\ -0.4323&-0.592&-0.23272863894943839798113773793091\\ -0.7303&0.04&-0.22959077803009316431117678104662\\ 0.1464&0.6149&0.13588682780065943976868637469759\\ -0.5154&0.0395&0.46102777383206591809202059407883\\ 0.6649&-0.1156&-0.22651115997910956793325851442687\\ 0.152&0.2539&-0.23985732806740791506457015735734\\ -0.03&0.0606&0.64396456614136038886542316026992\end{matrix}$ (3)
Figure 4: The intersection of $\hat{S}$ with the plane $y=0$, which is the boundary of two simply-connected components colored in yellow. The boxes contain the magnified view of their respective regions in the original slice.
We explain how we chose the triangulation $\mathcal{T}$ and obtained these coordinates. We start with all triangulations $\mathcal{T}_{n}=(V_{n},E_{n},F_{n})$ and polyhedral embeddings $S_{n}:(\Sigma_{2},\mathcal{T}_{n})\hookrightarrow\mathbb{R}^{3}$ from [HLZ07] for all 865 triangulations $n=1,\ldots,865$, where $S_{n}$ maps vertices of $\mathcal{T}_{n}$ to the integer lattice $\mathbb{Z}^{3}\cap[0,4]^{3}$. First, we apply a translation and scale the vectors so that the center-of-mass of the vertices is at the origin, and that $\max_{X\in V_{n}}\|X\|=\frac{1}{2}$. This ensures that $S_{n}(\Sigma_{2})\subset B_{1}(0)$, so we can treat $S_{n}$ as an embedding into the Klein model of $\mathbb{H}^{3}$. Then, for each $n$, we compute the hyperbolic angles of all triangles in $S_{n}$ and all the cone angles. We found that $S_{265}$ has the minimal cone angle difference to $2\pi$ among all other embeddings. After that, we apply a hill-climbing algorithm by randomly perturbing the coordinates of $S_{265}$ to minimize the objective function
$\mathcal{L}(S)=\max_{i=0,\ldots,9}|\theta_{i}(S)-2\pi|$
where $\theta_{i}(S)$ is the cone angle of $S$ at vertex $i$. This gives us a surface $S^{\prime}_{265}$ with $\mathcal{L}(S^{\prime}_{265})\leq 10^{-14}$. Let $z=(z_{0},\ldots,z_{9})$ denote the vector of $z$-coordinates of $S$. We implement a high-precision version of Newton’s method on the multivariable function
$\Theta(z)=\left(\theta_{0}(z)-2\pi,\theta_{1}(z)-2\pi,\ldots,\theta_{9}(z)-2\pi\right).$ (4)
We obtain the coordinates of $\hat{S}$ by truncating the $z$-coordinates at $10^{-32}$.
Figure 5: The intersection of $\hat{S}$ with the plane $x=0$, which is the boundary of a region that is not simply-connected. Again, the boxes contain the magnified view of their respective regions in the original slice.
1.3 A 12-Vertex Model with 2-Fold Symmetry
The candidate model $\hat{S}$ is rather unsatisfactory in the sense that it doesn’t admit any symmetry. On the other hand, the 8-vertex paper torus model from [Sch25] has a 2-fold symmetry that corresponds to a hyperelliptic involution, where the four fixed points are midpoints of 4 edges. Since every flat torus admits a hyperelliptic involution, Schwartz’s construction effectively reduced the dimension of the search space to find a family of paper torus in the hope of realizing every flat structure. Very recently, Doyle and Schwartz [DS25] found a family of 8-vertex flat polyhedral tori that realize any flat torus without reflection symmetry, which constitutes a dense subset in the moduli space of flat tori.
Figure 6: The embedding $\hat{S}_{12}$ with 2-fold rotational symmetry about the $z$-axis. Left: front view from the positive $x$-direction. Middle: bird’s-eye view from the positive $z$-direction. Right: view from the negative $z$-direction.
Motivated by this finding, I searched for a hyperbolic origami 2-torus with 2-fold symmetry that corresponds to a hyperelliptic involution. One first observes that if the order-2 symmetry acts freely on the set of vertices and fixes 6 points, then it must fix midpoints of 6 edges, which requires at least 12 vertices. I started by finding piecewise-linear embeddings of a 12-vertex triangulation of the 2-torus along with a candidate model with 2-fold symmetry into $\mathbb{R}^{3}$. A natural starting point is to consider the 7-regular triangulations, which has the largest symmetry group. I first checked whether the 7-regular triangulations of the 2-torus from [DU06] admit piecewise-linear embeddings into $\mathbb{R}^{3}$ with 2-fold symmetry. Then, I applied the same procedure as before to numerically approximate a hyperbolic origami 2-torus with 12 vertices and 2-fold symmetry. This led me to the following 12-vertex triangulation $\mathcal{T}$ with vertices indexed by $\{0,\ldots,11\}$. The indices of the 28 oriented triangles in $\mathcal{T}$ are listed below:
$\begin{matrix}(0,1,5)&(0,4,8)&(0,5,11)&(0,6,1)&(0,7,6)&(0,8,7)&(0,11,4)\\ (1,2,7)&(1,6,2)&(1,7,8)&(1,8,9)&(1,9,5)&(2,3,7)&(2,6,10)\\ (2,8,3)&(2,9,8)&(2,10,9)&(3,4,9)&(3,8,4)&(3,9,10)&(3,10,11)\\ (3,11,7)&(4,5,9)&(4,10,5)&(4,11,10)&(5,6,11)&(5,10,6)&(6,7,11)\end{matrix}$ (5)
Notice that $\mathcal{T}$ admits a $\mathbb{Z}/12\mathbb{Z}$ symmetry generated by the map $\sigma:i\mapsto(i+1)\mod 12$. The hyperelliptic involution is given by $\sigma^{6}$. Using this triangulation, I found a polyhedral embedding $\hat{S}_{12}:(\Sigma_{2},\mathcal{T})\hookrightarrow\mathbb{H}^{3}$ that satisfies $R\circ\hat{S}_{12}=\hat{S}_{12}$, where $R\in\operatorname{Isom}^{+}(\mathbb{H}^{3})$ is a rotation by $\pi$ about the $z$-axis. I checked that the cone angles of $\hat{S}_{12}$ are bounded away from $2\pi$ by $10^{-10}$. I provide the coordinates of $\hat{S}_{12}$ in $\S$5.3.
Figure 6 contains the visualization of the embedding $\hat{S}_{12}$ in the Klein model. It is very curious that when we use the Klein model to visualize $\hat{S}_{12}$, its appearance resembles a “double-deck pup tent” (the pup tent refers to the paper tori construction by Schwartz in [Sch25]). This hints at the possibility to generalize this construction.
Figure 7: The intersection of the 12-vertex model with the plane $x=0$ (left), $y=0$ (middle), and z=0.05z=-0.05 (right). One can clearly see the six midpoints of edges $(i,i+6)$ that are fixed by the rotational symmetry from the $x=0$ and $y=0$ slices.
The midpoints of the six edges $(i,i+6)$ all lie on the $z$-axis, which can be seen by slicing the embedding with planes that contains the $z$-axis. In Figure 7, we show the intersection of $\hat{S}_{12}$ with the planes $x=0$, $y=0$, and z=0.05z=-0.05. One can clearly see the six fixed points of the hyperelliptic involution from the $x=0$ and $y=0$ slices. Finally, in Figure 8, we show the induced triangle tiling of the universal cover $\mathbb{H}^{2}$ from $\hat{S}_{12}$ in the Poincaré disk model. In this visualization, the pairs of triangles $(i,j,k)$ and $(i+6,j+6,k+6)$ are colored the same way to reflect the 2-fold symmetry.
Figure 8: The isometrically induced 7-regular triangle tiling in the universal cover $\mathbb{H}^{2}$, visualized in the Poincaré disk model. The vertices $z_{0}$ and $z_{6}$ are fixed on the real axis, with $\Re(z_{0})<0$ and $z_{6}=-z_{0}$.
1.4 Paper Organization
This paper is organized as follows: In $\S$2, we remind the readers of the properties of the Klein model and explain how we incorporate floating-point arithmetic into the proofs. In $\S$3, we show that the cone angles of $\hat{S}$ differ from $2\pi$ by at most $10^{-28}$, and that $\hat{S}$ stays embedded after perturbing its $z$-coordinates by at most $10^{-7}$. In $\S$4, we analyze the Jacobian matrix of $\Theta(z)$ to show that there exists a hyperbolic origami 2-torus $S$ near $\hat{S}$.
An ongoing project is to generalize the 12-vertex construction to a family of hyperbolic origami 2-tori and check which hyperbolic structures can be realized.
1.5 Accompanying Programs
I have written programs in both Python and Mathematica to visualize the candidate models and slicing them with planes. Readers can access them via the following link:
1.6 Acknowledgements
This project is supported by an N.S.F. Research Grant DMS-2505281. I would like to thank my advisor, Richard Schwartz, for inspiring me to work on this problem. Much of the methodology used in this paper is inspired by [Sch25]. I would like to thank Peter Doyle, Fabian Lander, Alba Málaga, Stepan Paul, Steve Trettel, and Bena Tshishiku for helpful discussions on this topic. I would like to thank Frank Lutz for his manifolds webpage [Lut17] and Stefan Hougardy, Frank H. Lutz, Mariano Zelke for keeping a database of piecewise-affine embedded 2-tori with vertex-minimal triangulations, which helped me find prototypes for the candidate 2-torus in this paper.
2 Background
2.1 The Beltrami–Klein Model
Denote by $\mathbb{H}^{n}$ the hyperbolic $n$-space – the simply connected Riemannian $n$-manifold of constant sectional curvature -1. We mainly work with $\mathbb{H}^{3}$ in this paper. Here we provide the explicit construction known as the Beltrami–Klein model of $\mathbb{H}^{3}$. Consider the open unit Euclidean ball $B_{1}(0)\subset\mathbb{R}^{3}$. For each $X\in B_{1}(0)$, we write $a_{X}=1-\|X\|^{2}$. The Riemannian manifold $\mathbb{H}^{3}$ is defined as the manifold $B_{1}(0)$ along with the Riemannian metric $g$ given by
$g|_{X}=\left(\frac{(dx^{1})^{2}+(dx^{2})^{2}+(dx^{3})^{2}}{a_{X}}+\frac{(x^{1}dx^{1}+x^{2}dx^{2}+x^{3}dx^{3})^{2}}{a_{X}^{2}}\right)$ (6)
Let $V,W$ be two arbitrary vectors in $\mathbb{R}^{3}$, which we can identify with $T_{X}\mathbb{H}^{3}$. Then, $g|_{X}$ can be expressed as
$g|_{X}(V,W)=\frac{\langle V,W\rangle}{a_{X}}+\frac{\langle X,V\rangle\,\langle X,W\rangle}{a_{X}^{2}}=\frac{a_{X}\langle V,W\rangle+\langle X,V\rangle\langle X,W\rangle}{a_{X}^{2}}.$ (7)
where $\langle\cdot,\cdot\rangle$ denotes the Euclidean dot product. We will write $\langle V,W\rangle_{X}=g|_{X}(V,W)$ and $\|V\|_{X}=\sqrt{g|_{X}(V,V)}$ for the norm of $V\in T_{X}\mathbb{H}^{n}$.
Since geodesics in $\mathbb{H}^{3}$ are straight lines, a totally geodesic triangle in $\mathbb{H}^{3}$ with vertices $(X,Y,Z)$ is the same as the Euclidean triangle in the unit ball. We may compute the hyperbolic angle $\theta$ at $X$ as follows: Let V=YXV=Y-X and W=ZXW=Z-X. We may identify $V$ and $W$ with velocity vectors in $T_{X}\mathbb{H}^{3}$ of the geodesic from $X$ to $Y$ and $X$ to $Z$ respectively. Consider the following function:
$A(X,Y,Z)=\frac{\langle V,W\rangle_{X}^{2}}{\langle V,V\rangle_{X}\cdot\langle W,W\rangle_{X}}=\frac{(a_{X}\langle V,W\rangle+\langle X,V\rangle\langle X,W\rangle)^{2}}{(a_{X}\langle V,V\rangle+\langle X,V\rangle^{2})(a_{X}\langle W,W\rangle+\langle X,W\rangle^{2})}.$ (8)
Denote by $\sigma=\operatorname{sgn}(\langle V,W\rangle_{X})$ the sign of $\langle V,W\rangle_{X}$. It follows that
$\theta=\arccos\left(\frac{\langle V,W\rangle_{X}}{\|V\|_{X}\,\|W\|_{X}}\right)=\arccos\left(\sigma\sqrt{A(X,Y,Z)}\right);\hskip 10.0pt\cos\theta=\sigma\sqrt{A(X,Y,Z)}.$ (9)
2.2 Remark on Floating-Point Arithmetic
The proofs in this paper make use of computer floating-point arithmetic. To perform floating-point arithmetic up to high precision ($10^{-400}$ for our purposes), we use the Python package mpmath, which stores floating-point numbers of any designated precision and has built-in methods of basic arithmetic operations. When we claim that $\frac{a}{b}<\frac{c}{d}$ for floating-point numbers $a,b,c,d$, we implicitly carry out the computation of adbcad-bc in mpmath and check that the result is negative. We sometimes find equivalent fractions $\frac{a^{\prime}}{b^{\prime}}=\frac{a}{b}$ and $\frac{c^{\prime}}{d^{\prime}}=\frac{c}{d}$ with $a^{\prime},b^{\prime},c^{\prime},d^{\prime}\in\mathbb{Z}$ to simplify the computation into integer arithmetic.
For finding the bounds on the value of $e^{a}$ where $a$ is a floating-point number, see $\S$5.1.
3 The Candidate 2-Torus
In this section, we show that the cone angles of our candidate 2-torus $\hat{S}$ are very close to $2\pi$. We then show that if we apply perturbation to the $z$-coordinates of $\hat{S}$, the resulting 2-torus stays embedded as long as the perturbation is small enough.
Let $\mathcal{T}$ be a triangulation of the 2-torus $\Sigma_{2}$. A hyperbolic polyhedral 2-torus is an embedding $S:(\Sigma_{2},\mathcal{T})\hookrightarrow\mathbb{H}^{3}$ such that the image of the triangles in $\mathcal{T}$ under $S$ are geodesic triangles. A hyperbolic origami 2-torus is then a hyperbolic polyhedral 2-torus where the cone angle of each vertex is $2\pi$.
3.1 Robust Flatness
We start by defining flatness of a hyperbolic polyhedral 2-torus.
Definition 3.1.
We say that a hyperbolic polyhedral 2-torus $S$ is $\varepsilon$-flat if $\max_{i}|\theta_{i}-2\pi|<\varepsilon$, where $i\in\{0,\ldots,9\}$ and $\theta_{i}$ is the cone angle of vertex $i$.
Proposition 3.2.
$\hat{S}$ is $10^{-28}$-flat.
Proof.
Let $d_{i}$ be the degree of vertex $i$, and let $n_{0},\ldots,n_{d_{i}-1}$ be the indices of the neighbors of $i$, ordered so that $(i,n_{j},n_{j+1})\in F$ for $j=0,\ldots,d_{i}-1$ (here the subscripts $j$ and $j+1$ are taken modulo $d_{i}$). We first find integer vectors $Y_{i,n_{j}}\in\mathbb{R}^{2}$ for all $i=0,\ldots,9$ and $j=0,\ldots,d_{i}-1$ so that the difference $\delta_{i,j}=|\alpha_{i,j}-\beta_{i,j}|$ is minimized, where
$\alpha_{i,j}=A(\hat{X}_{i},\hat{X}_{n_{j}},\hat{X}_{n_{j+1}});\hskip 10.0pt\beta_{i,j}=\frac{\langle Y_{i,n_{j}},Y_{i,n_{j+1}}\rangle^{2}}{\langle Y_{i,n_{j}},Y_{i,n_{j}}\rangle\langle Y_{i,n_{j+1}},Y_{i,n_{j+1}}\rangle}$
Here, $A$ is the map from Equation (8). The coordinates of $Y_{i,n_{j}}$ are given in $\S$5.2. We obtained these coordinates by lifting the link of each vertex $\hat{X}_{i}$ isometrically into the Poincaré model of the universal cover $\mathbb{H}^{2}$, with $\hat{X}_{i}\mapsto 0$ and $\hat{X}_{n_{0}}$ mapping to the positive real axis. Then, we scale up each coordinate by $10^{32}$ to make them all integer-valued. We computed using floating-point arithmetic that $\max_{i,j}\delta_{i,j}\leq 2.93\times 10^{-32}$. We also computed that $\alpha_{i,j}\in[0.000052,0.918]$. It follows that $\alpha_{i,j},\beta_{i,j}\in[0.00005,0.92]$ for all $i,j$. In particular, $\langle X_{n_{j}}-X_{i},X_{n_{j+1}}-X_{i}\rangle_{X_{i}}$ and $\langle Y_{i,n_{j}},Y_{i,n_{j+1}}\rangle$ always share the same sign.
Next, we explain how we obtain the cone angle bounds from $\alpha_{i,j}$ and $\beta_{i,j}$. Consider two functions $\phi_{+}$ and $\phi_{-}$ given by $\phi_{+}(x)=\arccos(\sqrt{x})$ and $\phi_{-}(x)=\arccos(-\sqrt{x})$. Denote by $\theta_{i,j}$ the hyperbolic angle at $\hat{X}_{i}$ of the triangle $(\hat{X}_{i},X_{n_{j}},\hat{X}_{n_{j+1}})$, and denote by $\tau_{i,j}$ the Euclidean angle between the two vectors $Y_{i,n_{j}}$ and $Y_{i,n_{j+1}}$. Since $\langle X_{n_{j}}-X_{i},X_{n_{j+1}}-X_{i}\rangle_{X_{i}}$ and $\langle Y_{i,n_{j}},Y_{i,n_{j+1}}\rangle$ always share the same sign, we know that
  • $\theta_{i,j}=\phi_{+}(\alpha_{i,j})$ and $\tau_{i,j}=\phi_{+}(\beta_{i,j})$ if $\langle X_{n_{j}}-X_{i},X_{n_{j+1}}-X_{i}\rangle_{X_{i}}>0$;
  • $\theta_{i,j}=\phi_{-}(\alpha_{i,j})$ and $\tau_{i,j}=\phi_{-}(\beta_{i,j})$ if $\langle X_{n_{j}}-X_{i},X_{n_{j+1}}-X_{i}\rangle_{X_{i}}<0$.
Next, observe that
$\frac{d\phi_{+}}{dx}=\frac{-1}{2\sqrt{x(1-x)}};\hskip 10.0pt\frac{d\phi_{-}}{dx}=\frac{1}{2\sqrt{x(1-x)}}$
One can compute directly that $|\frac{d\phi_{+}}{dx}|,|\frac{d\phi_{-}}{dx}|\leq 71$ on $[0.00005,0.92]$, so both $\phi_{+}$ and $\phi_{-}$ are $71$-Lipschitz on $[0.00005,0.92]$. Finally, we have
$|\theta_{i}-2\pi|$ $=\left|\sum_{j=0}^{d_{i}-1}\theta_{i,j}-\sum_{j=0}^{d_{i}-1}\tau_{i,j}\right|\leq\sum_{j=0}^{d_{i}-1}|\theta_{i,j}-\tau_{i,j}|\leq\max_{i}d_{i}\cdot\max_{i,j}|\theta_{i,j}-\tau_{i,j}|$
[…]
Taking $\varepsilon=10^{-28}$ completes the proof. ∎
3.2 Robust Embeddedness
The next step is to show that if we slightly perturb the $z$-coordinates of $\hat{S}$, the resulting 2-torus is still embedded. The key observation (and the main reason we use the Beltrami–Klein model of $\mathbb{H}^{3}$) is that the underlying set of the hyperbolic geodesic triangle with vertices $V_{1},V_{2},V_{3}\in\mathbb{H}^{3}$ is precisely the Euclidean triangle with the same vertices. Therefore, in this section we “forget” the hyperbolic structure and instead treat $\hat{S}$ as a map $\hat{S}:(\Sigma_{2},\mathcal{T})\rightarrow B_{1}(0)$ whose image is the unit Euclidean ball. We begin with the following definition, which is not “natural” in the hyperbolic setting but is more appropriate for Euclidean geometry.
Definition 3.3.
A map $S:(\Sigma_{2},\mathcal{T})\rightarrow B_{1}(0)$ is $\lambda$-robustly embedded if for all $S^{\prime}$ obtained from moving each of the $z$-coordinates of the vertices of $S$ by at most $\lambda$ is also embedded.
Remark 3.4.
Readers may have noticed that Definition 3.3 is not invariant under hyperbolic isometries. That is, suppose $S:(\Sigma_{2},\mathcal{T})\hookrightarrow\mathbb{H}^{3}$ is $\lambda$-robustly embedded and $M\in\operatorname{Isom}(\mathbb{H}^{3})$, then there is no reason whatsoever to expect that $M\circ S$ must also be $\lambda$-robustly embedded. One should keep in mind that throughout this section, we will “pretend to” work with Euclidean geometry.
To show embeddedness, it suffices to consider pairs of triangles with no common vertices and pairs that share one vertex in common. Below is a lemma from [Sch25]. We present a self-contained proof here.
Lemma 3.5 ([Sch25, Lemma 3.2]).
Given a map $S:(\Sigma_{2},\mathcal{T})\rightarrow B_{1}(0)$ sending triangles in $\mathcal{T}$ to Euclidean triangles in $B_{1}(0)$. If all pairs of triangles of $S$ sharing no vertices in common are disjoint and all pairs of triangles sharing one vertex only intersect at the shared vertex, then $S$ is embedded.
Proof.
Consider the pairs of triangles $(i,j,k)$ and $(j,i,l)$ that share two vertices. Denote by $e_{ab}$ the edge connecting $X_{a}$ to $X_{b}$, where $a,b\in\{i,j,k,l\}$. We first recall that every edge is contained in two triangles, so the edge $e_{ik}$ is contained in another triangle $(i,k,m)$, where $m\neq j$. If the interior of $e_{ik}$ intersects $(j,i,l)$, then $(j,i,l)$ and $(i,k,m)$ intersect at a point other than $X_{i}$. Notice that $l\neq m$, for otherwise the degree of vertex $i$ would equal 3. This implies $(j,i,l)$ and $(i,k,m)$ only share one vertex, but that contradicts our assumption in the lemma. The same argument shows the interior of $e_{jk}$ doesn’t intersect $(j,i,l)$, and the interior of $e_{il}$ and $e_{jl}$ doesn’t intersect $(i,j,k)$, so these two triangles only intersect at $e_{ij}$, which shows that $S$ is embedded. ∎
We first deal with the pairs of disjoint triangles. We say that two disjoint triangles $T_{1}$ and $T_{2}$ in $\mathbb{R}^{3}$ are $\delta$-separated if they remain disjoint after simultaneously perturbing the $z$-coordinates of the vertices by at most $\delta$. Here is a sufficient hyperplane separation test to show that $T_{1}$ and $T_{2}$ are $\delta$-separated.
Lemma 3.6.
Two disjoint triangles $T_{1}$ and $T_{2}$ are $\delta$-separated if there exists a vector $\hat{N}$ with $\|\hat{N}\|_{\infty}<C$ for some $C\in\mathbb{R}$, such that
$\min_{X\in T_{1}}\langle X,\hat{N}\rangle\,-\,\max_{Y\in T_{2}}\langle Y,\hat{N}\rangle\,>\,2\delta C.$ (10)
In particular, if $T_{1}$ has vertices $V_{1},V_{2},V_{3}$ and $T_{2}$ has vertices $W_{1},W_{2},W_{3}$, then it suffices to check that
$\min_{X\in\{V_{1},V_{2},V_{3}\}}\langle X,\hat{N}\rangle\,-\,\max_{Y\in\{W_{1},W_{2},W_{3}\}}\langle Y,\hat{N}\rangle\,>\,2\delta C.$ (11)
Proof.
We know that every $X\in T$ can be expressed as $X=t_{1}V_{1}+t_{2}V_{2}+t_{3}V_{3}$ where $t_{i}\geq 0$ and $t_{1}+t_{2}+t_{3}=1$. Linearity of the dot product then implies the minimum of $\langle X,\hat{N}\rangle$ must be achieved at one of the vertices of $T$, so Equation (11) implies (10). Now, if $T^{\prime}_{1}$ and $T^{\prime}_{2}$ are obtained from purturbing the $z$-coordinates of the vertices of $T_{1}$ and $T_{2}$ by at most $\delta$, then the dot product $\langle V_{i},\hat{N}\rangle$ is perturbed by at most $\delta C$. It follows that
$\min_{X^{\prime}\in T_{1}^{\prime}}\langle X^{\prime},\hat{N}\rangle-\max_{Y^{\prime}\in T_{2}^{\prime}}\langle Y^{\prime},\hat{N}\rangle\geq\min_{X\in T_{1}}\langle X,\hat{N}\rangle-\max_{Y\in T_{2}}\langle Y,\hat{N}\rangle-2\delta C>0.$
This implies $T_{1}^{\prime}\cap T_{2}^{\prime}=\emptyset$, which completes the proof. ∎
Next, we deal with triangles that share one of their vertices. Suppose $(U,V_{1},V_{2},W_{1},W_{2})$ are five points in $\mathbb{R}^{3}$, and $T_{1}$, $T_{2}$ are two triangles with vertices $(U,V_{1},V_{2})$ and $(U,W_{1},W_{2})$ respectively. We say that $T_{1}$ and $T_{2}$ are $\delta$-separated if for all $T^{\prime}_{1}$ and $T^{\prime}_{2}$ with vertices $(U^{\prime},V^{\prime}_{1},V^{\prime}_{2})$ and $(U^{\prime},W^{\prime}_{1},W^{\prime}_{2})$ such that $(U^{\prime},V^{\prime}_{1},V^{\prime}_{2},W^{\prime}_{1},W^{\prime}_{2})$ are obtained from purturbing the $z$-coordinates of $(U,V_{1},V_{2},W_{1},W_{2})$ by at most $\delta$, $T^{\prime}_{1}\cap T^{\prime}_{2}=\{U^{\prime}\}$.
Lemma 3.7.
$T_{1}$ and $T_{2}$ are $\delta$-separated if there exists a vector $\hat{N}$ with $\|\hat{N}\|_{\infty}<C$ for some $C\in\mathbb{R}$, such that
$\min_{X\in\{V_{1},V_{2}\}}\langle X,\hat{N}\rangle-\langle U,\hat{N}\rangle>2\delta C\hskip 5.0pt\text{and}\hskip 5.0pt\langle U,\hat{N}\rangle-\min_{Y\in\{W_{1},W_{2}\}}\langle Y,\hat{N}\rangle>2\delta C.$ (12)
Proof.
Suppose $T^{\prime}_{1}$ and $T^{\prime}_{2}$ are obtained from purturbing $(U,V_{1},V_{2},W_{1},W_{2})$ by at most $\delta$. The dot products with $\hat{N}$ is perturbed by at most $\delta C$. It follows that
$\min_{X^{\prime}\in\{V^{\prime}_{1},V^{\prime}_{2}\}}\langle X^{\prime},\hat{N}\rangle-\langle U^{\prime},\hat{N}\rangle>0\hskip 5.0pt\text{and}\hskip 5.0pt\langle U^{\prime},\hat{N}\rangle-\min_{Y^{\prime}\in\{W^{\prime}_{1},W^{\prime}_{2}\}}\langle Y^{\prime},\hat{N}\rangle>0.$ (13)
To see this implies $T^{\prime}_{1}\cap T^{\prime}_{2}=\{U^{\prime}\}$, suppose there exists $X^{\prime}\in T^{\prime}_{1}\cap T^{\prime}_{2}$. We may express $X^{\prime}=t_{0}U+t_{1}V^{\prime}_{1}+t_{2}V^{\prime}_{2}=s_{0}U+s_{1}W^{\prime}_{1}+s_{2}W^{\prime}_{2}$ where $t_{0}+t_{1}+t_{2}=s_{0}+s_{1}+s_{2}=1$ and all coefficients $t_{i},s_{j}\geq 0$. It follows that
$\langle X^{\prime},\hat{N}\rangle-\langle U^{\prime},\hat{N}\rangle=t_{1}\langle V^{\prime}_{1}-U^{\prime},\hat{N}\rangle+t_{2}\langle V^{\prime}_{2}-U^{\prime},\hat{N}\rangle$
and
$\langle U^{\prime},\hat{N}\rangle-\langle X^{\prime},\hat{N}\rangle=s_{1}\langle U^{\prime}-W^{\prime}_{1},\hat{N}\rangle+s_{2}\langle U^{\prime}-W^{\prime}_{2},\hat{N}\rangle$.
This implies
$t_{1}\langle V^{\prime}_{1}-U^{\prime},\hat{N}\rangle+t_{2}\langle V^{\prime}_{2}-U^{\prime},\hat{N}\rangle+s_{1}\langle U^{\prime}-W^{\prime}_{1},\hat{N}\rangle+s_{2}\langle U^{\prime}-W^{\prime}_{2},\hat{N}\rangle=0,$
so Equation (13) implies $t_{1}=t_{2}=s_{1}=s_{2}=0$, which gives us $X^{\prime}=U^{\prime}$ as desired. ∎
We are now ready to prove that $\hat{S}$ is robustly embedded.
Proposition 3.8.
$\hat{S}$ is $10^{-7}$-robustly embedded.
Proof.
We dilate the surface $\hat{S}$ by a factor of $10^{32}$ so that every vertex has integer coordinates. Call the dilated surface $\hat{S}^{\prime}$. Lemma 3.6 and 3.7 tells us that it suffices to find normal vectors $\hat{N}$ for each pair of triangles in $\hat{S}^{\prime}$.
Among the 276 pairs of triangles from $F$, there are 82 pairs that share no vertices in common, and 158 of which share one vertex in common. We provide an algorithm to generate $\hat{N}$ for each of these pairs where the conditions in Lemma 3.6 and 3.7 hold. Consider a map $\rho:\mathbb{Z}_{>0}\rightarrow\mathbb{Z}^{3}$ as follows:
$\rho(n)=\left(\lfloor 10^{5}\cdot L(n\sqrt{2})\rfloor,\lfloor 10^{5}\cdot L(n\sqrt{3})\rfloor,\lfloor 10^{5}\cdot L(n\sqrt{5})\rfloor\right);\hskip 10.0ptL(x)=2(x-\lfloor x\rfloor)-1$
It is clear that $\|\rho(n)\|_{\infty}<10^{5}$ for all $n$. We found indices $n<2000$ such that the conditions in Lemma 3.6 hold for $\delta=10^{25}$ by taking $\hat{N}=\pm\rho(n)$ for all but the following two pairs of disjoint triangles, where we manually found the candidates for $\hat{N}$:
[…]
We also computed $\rho(n)$ for the triangle pairs that share one vertex. We found indices $n<10^{5}$ such that the conditions in Lemma 3.7 hold for $\delta=10^{25}$ by taking $\hat{N}=\pm\rho(n)$ for all but the following nine pairs, where we manually found the candidates for $\hat{N}$:
[…]
Lemma 3.5 then implies the $10^{32}$-dilated surface $\hat{S}^{\prime}$ is $10^{25}$-robustly embedded, so the original surface $\hat{S}$ is $10^{7}$-robustly embedded, as desired. ∎
4 Finding Origami 2-Tori
Our method of finding the flat 2-tori is analogous to Schwartz’s method in [Sch25]. The idea is to use the inverse function theorem to show that there exists a neighborhood around $\hat{z}$ whose image under the map $\Theta$ in Equation (4) hits the origin.
Definition 4.1.
Let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be a smooth map. We say that $f$ is $\lambda$-expansive on a subset $A\subset\mathbb{R}^{n}$ if for all unit vectors $V\in\mathbb{R}^{n}$ and $p,q\in A$, we have
  • $\|df|_{p}(V)\|>2\lambda$;
  • The angle between $df|_{p}(V)$ and $df|_{q}(V)$ is less than $\frac{\pi}{3}$.
The following lemma is proved by Schwartz in [Sch25]. Here we provide an alternative self-contained proof.
Lemma 4.2 ([Sch25, Lemma 3.3]).
Let $B_{r}(p)\subset\mathbb{R}^{n}$ denote the open ball of radius $r$ centered at $p$. If $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is $\lambda$-expansive on $B_{r}(p)$ for some $\lambda>0$, then $B_{\lambda r}(f(p))\subset f(B_{r}(p))$.
Proof.
After normalizing by translation and scaling we may assume $f(p)=p=0$, $\lambda=r=1$. Since $\|df|_{q}(V)\|>0$ for all $q\in B_{1}(0)$ and unit vectors $V$, $df|_{q}$ is nondegenerate, so by the inverse function theorem, $f$ is an open map on $B_{1}(0)$. It follows that $\mathbb{R}^{n}\setminus f(B_{1}(0))$ is closed. Consider the set $A=\overline{B_{s}(0)}\setminus f(B_{1}(0))$, where $s=2\sup_{x\in B_{1}(0)}\|f(x)\|$. We know that $A$ is compact, and $A\cup f(B_{1}(0))=\overline{B_{s}(0)}$. The statement is true if $\|x\|\geq 1$ for all $x\in A$.
[…]
Proposition 4.3.
[…]
Proof of Theorem 1.1 assuming Proposition 4.3.
[…]
The rest of this section will be devoted to proving Proposition 4.3.
4.1 The Jacobian
Our first goal is to analytically compute the Jacobian matrix of the map $\Theta$. Recall from Equation (4) that $\Theta$ is given by
$\Theta_{i}(z)=\sum_{(i,j,k)\in F}\theta_{ijk}(z)-2\pi,$
where $\theta_{ijk}(z)$ denotes the angle of the triangle $(i,j,k)$ at vertex $i$.
We start by computing the first-order partials of $\theta_{ijk}$. Given $X_{i}=(x_{i},y_{i},z_{i})$, denote by $\partial_{i}$ the operator $\frac{\partial}{\partial z_{i}}$. Consider a triangle $(i,j,k)\in F$. We again write $V=X_{j}-X_{i}$, $W=X_{k}-X_{i}$, and $a_{X_{i}}=1-\|X_{i}\|^{2}$. We would like to compute the action of $\partial_{i},\partial_{j},\partial_{k}$ on $\langle V,W\rangle_{X_{i}}$, $\|V\|_{X_{i}}$, and $\|W\|_{X_{i}}$. By symmetry of the Riemannian metric, it suffices to compute $\partial_{i}$ and $\partial_{j}$. We first observe that
$\begin{matrix}\partial_{i}\langle X_{i},X_{i}\rangle=2z_{i}&\partial_{i}\langle V,V\rangle=2z_{i}-2z_{j}&\partial_{i}\langle W,W\rangle=2z_{i}-2z_{k}\\ \partial_{i}\langle X_{i},V\rangle=z_{j}-2z_{i}&\partial_{i}\langle X_{i},W\rangle=z_{k}-2z_{i}&\partial_{i}\langle V,W\rangle=2z_{i}-z_{j}-z_{k}\\ \partial_{j}\langle V,X_{i}\rangle=z_{i}&\partial_{j}\langle V,W\rangle=z_{k}-z_{i}&\partial_{j}\langle V,V\rangle=2z_{j}-2z_{i}\end{matrix}$
It follows that
$\partial_{i}\langle V,W\rangle_{X_{i}}=\partial_{i}\left(\frac{\langle V,W\rangle}{a_{X_{i}}}+\frac{\langle X_{i},V\rangle\langle X_{i},W\rangle}{a_{X_{i}}^{2}}\right)$ (14)
$=$ $\frac{2z_{i}-z_{j}-z_{k}}{a_{X_{i}}}+\frac{2z_{i}\langle V,W\rangle+(z_{j}-2z_{i})\langle X_{i},W\rangle+(z_{k}-2z_{i})\langle X_{i},V\rangle}{a_{X_{i}}^{2}}+\frac{4z_{i}\langle X_{i},V\rangle\langle X_{i},W\rangle}{a_{X_{i}}^{3}}$
and
$\partial_{j}\langle V,W\rangle_{X_{i}}=\partial_{j}\left(\frac{\langle V,W\rangle}{a_{X_{i}}}+\frac{\langle X_{i},V\rangle\langle X_{i},W\rangle}{a_{X_{i}}^{2}}\right)=\frac{z_{k}-z_{i}}{a_{X_{i}}}+\frac{z_{i}\langle X_{i},W\rangle}{a_{X_{i}}^{2}}.$ (15)
We also have
$\partial_{i}\|V\|_{X_{i}}$ $=\partial_{i}\sqrt{\langle V,V\rangle_{X_{i}}}=\frac{1}{2\|V\|_{X_{i}}}\partial_{i}\langle V,V\rangle_{X_{i}}$ (16)
$=\frac{1}{\|V\|_{X_{i}}}\left(\frac{z_{i}-z_{j}}{a_{X_{i}}}+\frac{z_{i}\langle V,V\rangle+(z_{j}-2z_{i})\langle X_{i},V\rangle}{a_{X_{i}}^{2}}+\frac{2z_{i}\langle X_{i},V\rangle^{2}}{a_{X_{i}}^{3}}\right)$
and
$\partial_{j}\|V\|_{X_{i}}=\frac{1}{2\|V\|_{X_{i}}}\partial_{j}\langle V,V\rangle_{X_{i}}=\frac{1}{\|V\|_{X_{i}}}\left(\frac{z_{j}-z_{i}}{a_{X_{i}}}+\frac{z_{i}\langle X_{i},V\rangle}{a_{X_{i}}^{2}}\right).$ (17)
Therefore, for all $l\in\{i,j,k\}$, we have
$\partial_{l}\theta_{ijk}$ $=\partial_{l}\arccos\left(\frac{\langle V,W\rangle_{X_{i}}}{\|V\|_{X_{i}}\|W\|_{X_{i}}}\right)=\frac{-\|V\|_{X_{i}}\|W\|_{X_{i}}}{\sqrt{\langle V,V\rangle_{X_{i}}\langle W,W\rangle_{X_{i}}-\langle V,W\rangle_{X_{i}}^{2}}}\cdot\partial_{l}\left(\frac{\langle V,W\rangle_{X_{i}}}{\|V\|_{X_{i}}\|W\|_{X_{i}}}\right)$ (18)
$=\frac{\partial_{l}(\|V\|_{X_{i}}\|W\|_{X_{i}})\langle V,W\rangle_{X_{i}}}{\|V\|_{X_{i}}\|W\|_{X_{i}}\cdot\sqrt{\langle V,V\rangle_{X_{i}}\langle W,W\rangle_{X_{i}}-\langle V,W\rangle_{X_{i}}^{2}}}-\frac{\partial_{l}\langle V,W\rangle_{X_{i}}}{\sqrt{\langle V,V\rangle_{X_{i}}\langle W,W\rangle_{X_{i}}-\langle V,W\rangle_{X_{i}}^{2}}}.$
Next, consider the following 10-by-10 matrix $M$:
$\begin{pmatrix}-7.526&0.793&-0.228&-0.264&-0.782&1.372&1.173&-0.098&0.192&2.158\\ 0.684&4.991&-0.344&-0.104&-1.468&-0.260&0.208&-0.661&-1.528&0\\ -0.203&-0.356&3.913&-1.008&-0.104&0&-0.091&-0.588&-0.017&0.224\\ -0.255&-0.116&-1.093&4.665&-0.847&-0.035&0.117&-0.133&-0.037&0\\ -0.753&-1.639&-0.112&-0.843&5.270&-0.642&0.820&0.015&0&0.870\\ 1.112&-0.244&0&-0.029&-0.540&-2.245&0.162&0&-0.369&1.354\\ 1.006&0.207&-0.087&0.104&0.730&0.172&-9.148&0&0&3.639\\ -0.086&-0.674&-0.580&-0.121&0.013&0&0&5.063&-1.681&0\\ 0.128&-1.185&-0.013&-0.026&0&-0.305&0&-1.277&3.455&0\\ 1.751&0&0.204&0&0.733&1.355&3.444&0&0&-11.599\end{pmatrix}$
We state the following lemma, which we will prove in the next section.
Lemma 4.4.
[…]
$\left|\frac{\partial^{2}\Theta_{i}}{\partial z_{j}\partial z_{k}}(z)\right|\leq 10^{14}.$
Corollary 4.5.
[…]
$d\Theta|_{z}=M+E,$
where $M$ is the matrix above and $E$ is a matrix with $\|E\|_{\infty}<0.002$.
Proof.
Let $r=10^{-18}$. One can use high precision floating-point arithmetic to compute $d\Theta|_{\hat{z}}$ and check that $d\Theta|_{\hat{z}}=M+E^{\prime}$, where $\|E^{\prime}\|_{\infty}<0.001$. It then suffices to show that $d\Theta|_{z}=d\Theta|_{\hat{z}}+E^{\prime\prime}$, where $\|E^{\prime\prime}\|_{\infty}<0.001$. To see this, notice that $\|z-\hat{z}\|<r$. It follows that
$\left|\frac{\partial\Theta_{i}}{\partial z_{j}}(z)-\frac{\partial\Theta_{i}}{\partial z_{j}}(\hat{z})\right|\leq 10\cdot r\cdot\sup_{k\in[10],w\in B}\left|\frac{\partial^{2}\Theta_{i}}{\partial z_{j}\partial z_{k}}(w)\right|<10^{-3}.$
This finishes the proof. ∎
Proof of Proposition 4.3.
We first claim that $\|M(V)\|>1.5$ for all unit vectors $V$. To see this, one can numerically show that the smallest root of the characteristic polynomial of $M^{t}M$ is greater than $2.25$, where $M^{t}$ denotes the transpose of $M$. It then follows that the smallest singular value of $M$ is greater than $1.5$ (see Remark 4.6).
Next, given a 10-by-10 matrix $E$ with $\|E\|_{\infty}<0.002$, we must have
$\|E\|_{2}\leq\|E\|_{F}\leq 10^{2}\|E\|_{\infty}<0.2.$
Here, $\|E\|_{2}=\sup_{\|V\|=1}\|E(V)\|$ is the spectral norm of $E$ and $\|E\|_{F}$ is the Frobenius norm. It follows that
$\|d\Theta|_{z}(V)\|=\|M(V)+E(V)\|\geq|\|M(V)\|-\|E(V)\||\geq 1.5-0.2>1.$
This shows that $\|d\Theta|_{z}(V)\|>2\lambda$.
[…]
$\sin\theta\leq\frac{\|E^{\prime}(V)\|}{\|d\Theta|_{z}(V)\|}\leq\frac{0.4}{1.5}<\frac{\sqrt{3}}{2}.$
This shows that $\theta<\frac{\pi}{3}$. ∎
Remark 4.6.
We explain how we find that the singular values of $M$ are greater than $1.5$. Since the entries of $M^{t}M$ are obtained from taking sums and products of entries of $M$, we may compute them precisely using floating-point arithmetic. This further implies that we may compute precisely the coefficients of the characteristic polynomial $\chi$. We find that the characteristic polynomial has 10 distinct roots $x_{i}$ by finding $a_{i},b_{i}\in\mathbb{R}$ such that $a_{i}<b_{i}$ and $\chi(a_{i})\chi(b_{i})<0$. The intermediate value theorem then guarantees the existence of the root $x_{i}\in(a_{i},b_{i})$ for each of the ten pairs $(a_{i},b_{i})$. We omit the tedious computation.
4.2 Bounds on Angles and Hyperbolic Distances
Before proving Lemma 4.4, we need to first establish some bounds on the angles and hyperbolic distances of the embedded surfaces in $\mathcal{B}$.
Lemma 4.7.
Consider the closed ball $B_{r}(0)\subset\mathbb{R}^{3}$ of radius $r\in(0,1)$ centered at the origin. Then, for all $X\in B_{r}(0)$, $V\in\mathbb{R}^{3}$, we have
$\|V\|^{2}\leq\|V\|_{X}^{2}\leq\frac{\|V\|^{2}}{(1-r^{2})^{2}}.$ (19)
Proof.
The first inequality follows from monotonicity of $1/(1+x^{2})$ n $[0,r]$:
$\langle V,V\rangle_{X}=\frac{\langle V,V\rangle}{a_{X}}+\frac{\langle V,X\rangle^{2}}{a_{X}^{2}}\geq\frac{\|V\|^{2}}{a_{X}}\geq\|V\|^{2}.$
The second inequality follows from a routine application of the Cauchy-Schwarz inequality and monotonicity of $1/(1-x^{2})^{2}$ on $[0,r]$:
$\langle V,V\rangle_{X}\leq\frac{\|V\|^{2}}{a_{X}}+\frac{\|V\|^{2}\|X\|^{2}}{a_{X}^{2}}=\frac{\|V\|^{2}}{a_{X}^{2}}\leq\frac{\|V\|^{2}}{(1-r^{2})^{2}}.$
This completes the proof. ∎
Corollary 4.8.
Given $S\in\mathcal{B}$ and an edge $e$ connecting $X$ to $Y$, let V=YXV=Y-X. Then, we have $\|V\|_{X}\in[0.5,13]$.
Proof.
Denote by $\hat{V}=\hat{Y}-\hat{X}$, where $\hat{V}$ is an edge connecting $\hat{X}$ to $\hat{Y}$ in $\hat{S}$ that corresponds to $V$. We compute numerically using Equation (6) that $\|\hat{V}\|\in[0.509,1.561]$. Then, since $\hat{S}\subset B_{0.79}(0)$, we must have $S\subset B_{0.8}(0)$ for all $S\in\mathcal{B}$, so applying Lemma 4.7 with $r=0.8$ gives us the desired bounds. ∎
Proposition 4.9.
Given $S\in\mathcal{B}$, let $e$ be an arbitrary edge in $S$ connecting $X$ to $Y$ that corresponds to the edge $\hat{e}$ in $\hat{S}$ connecting $\hat{X}$ to $\hat{Y}$. Denote by $l(e)$ and l(e^)l($\hat${e}) the hyperbolic length of edges $e$ and $\hat{e}$ respectively. Then, we have
$|l(e)-l(\hat{e})|\leq 1.6\times 10^{-17}.$
Proof.
[…]
$d_{g}(X,\hat{X})=l(\gamma(t))=\int_{0}^{1}\|\gamma^{\prime}\|_{\gamma(t)}dt=\int_{0}^{1}\|\hat{X}-X\|_{\gamma(t)}dt\leq 8\|\hat{X}-X\|\leq 8\times 10^{-18}$
where the second to the last inequality comes from Lemma 4.7 and $1/(1-0.8^{2})^{2}\leq 8$. We similarly get $d_{g}(Y,\hat{Y})\leq 8\times 10^{-18}$. It follows that
[…]
Combining these two equations gives us the desired result. ∎
Corollary 4.10.
With the same setup as above, we have $l(e)\in[0.6,2.1]$.
Proof.
Using the coordinates of $\hat{S}$, we compute that $l(\hat{e})\in[0.63,2.08]$. We explain how we establish the bound. Given two distinct points $X,Y\in\mathbb{H}^{3}$, the quadratic polynomial $\|(1-t)X+tY\|^{2}=1$ has two distinct real roots $t_{1}<0$ and $t_{2}>1$. Let $A=(1-t_{1})X+t_{1}Y$ and $B=(1-t_{2})X+t_{2}Y$. The hyperbolic distance $d_{g}(X,Y)$ is given by
$d_{g}(X,Y)=\frac{1}{2}\ln\left(\frac{\|X-B\|\|Y-A\|}{\|X-A\|\|Y-B\|}\right)=\frac{1}{2}\ln\left(\frac{t_{2}-t_{1}t_{2}}{t_{1}-t_{1}t_{2}}\right).$
We may express $\|(1-t)X+tY\|^{2}=1$ as $at^{2}+bt+c=0$, where
$a=\langle Y-X,Y-X\rangle;\hskip 4.0ptb=2\langle X,Y-X\rangle;\hskip 4.0ptc=\langle X,X\rangle-1.$
Denote by $\Delta=b^{2}-4ac$ the discriminant of the quadratic $at^{2}+bt+c=0$. One can check that $\Delta>0$ if $X\neq Y$. It follows that
$d_{g}(X,Y)=\frac{1}{2}\ln\left(\frac{\Delta^{1/2}-b-2c}{-\Delta^{1/2}-b-2c}\right)=\ln\left(\Delta^{1/2}-b-2c\right)-\frac{1}{2}\ln\left(4c^{2}+4ac+4bc\right)$ (20)
Since the square root function is monotonic increasing, it suffices to find floating-point numbers $x_{1},x_{2}$ such that $x_{1}^{2}\leq x\leq x_{2}^{2}$ to show $\sqrt{x}\in[x_{1},x_{2}]$. This allows us to bound the value of $\Delta^{1/2}$ within $10^{-32}$. It follows that $\Delta^{1/2}-b-2c\in[1.93,6.3]$. We also directly computed that $4c^{2}+4ac+4bc\in[0.62,1.99]$. It remains to compute $\ln(x)$ for $x\in[0.62,6.3]$, which gives us $l(\hat{e})\in[0.63,2.08]$. We refer to $\S$5.1 for how we obtained this bound. The corollary then follows from Proposition 4.9. ∎
Proposition 4.11.
The map
$\psi(a,b,c)=\frac{\cosh a\cosh b-\cosh c}{\sinh a\sinh b}$ (21)
is $70$-Lipschitz on $0.6,2.1]^{3}$.
Proof.
The partial derivatives of $\psi$ are given by
$\frac{\partial\psi}{\partial a}=\frac{\cosh a\cosh c-\cosh b}{\sinh^{2}a\sinh b};\hskip 10.0pt\frac{\partial\psi}{\partial b}=\frac{\cosh b\cosh c-\cosh a}{\sinh^{2}b\sinh a};\hskip 10.0pt\frac{\partial\psi}{\partial c}=-\frac{\sinh c}{\sinh a\sinh b}.$
Since $\sinh,\cosh,\tanh$ are all positive and monotonic increasing on $[0.5,2.1]$, we have $|\frac{\partial\psi}{\partial c}|\leq\frac{\sinh 2.1}{\sinh^{2}0.5}\leq 15$ and
$\left|\frac{\partial\psi}{\partial a}\right|\leq\left|\frac{\cosh c}{\tanh a\sinh a\sinh b}\right|+\left|\frac{1}{\sinh^{2}a\tanh b}\right|\leq\frac{\cosh 2.1+1}{\tanh 0.5\sinh^{2}0.5}\leq 42.$ (22)
See $\S$5.1 for finding bounds on the values of $\sinh x,\cosh x,\tanh x$ where $x=0.5,2.1$. Similarly, $|\frac{\partial\psi}{\partial b}|\leq 42$. This implies
$\frac{|\psi(y)-\psi(x)|}{\|y-x\|}\leq\sup_{I}\sqrt{\left(\frac{\partial\psi}{\partial a}\right)^{2}+\left(\frac{\partial\psi}{\partial b}\right)^{2}+\left(\frac{\partial\psi}{\partial c}\right)^{2}}\leq\sqrt{42^{2}+42^{2}+15^{2}}\leq 70$
for all $x,y\in[0.6,2.1]^{3}$. ∎
Corollary 4.12.
For all $S\in\mathcal{B}$ and an angle $\theta$ of a triangle in $S$, we have $|\sin\theta|\geq 0.24$.
Proof.
Let $\hat{\theta}$ be the corresponding angle of $\theta$ in $\hat{S}$. We compute using Equation (8) and (9) that $\hat{\theta}$ satisfies $\cos\hat{\theta}\in[-0.008,0.96]$. Denote by $a,b,c$ the length of the edges of the triangle that contains $\theta$, where $c$ corresponds to the opposite edge of $\theta$. Denote by $\hat{a},\hat{b},\hat{c}$ the edge lengths in $\hat{S}$. From Proposition 4.9, we know that $|a-\hat{a}|\leq 1.6\times 10^{-17}$. Similarly, $|b-\hat{b}|\leq 1.6\times 10^{-17}$ and $|c-\hat{c}|\leq 1.6\times 10^{-17}$. It follows that $|\sqrt{a^{2}+b^{2}+c^{2}}-\sqrt{\hat{a}^{2}+\hat{b}^{2}+\hat{c}^{2}}|\leq 2.8\times 10^{-17}$. Also, the hyperbolic law of cosine implies
$\cos\theta=\frac{\cosh a\cosh b-\cosh c}{\sinh a\sinh b}=\psi(a,b,c).$
From Proposition 4.11 we know that $\psi$ is $70$-Lipschitz on $0.6,2.1]^{3}$. Also, Proposition 4.10 gives us $a,b,c,\hat{a},\hat{b},\hat{c}\in[0.6,2.1]$, so $|\cos\theta-\cos\hat{\theta}\,|\leq 2\times 10^{-15}$. This gives us $\cos\theta\in[-0.01,0.961]$. It remains a routine computation to check that $|\sin\theta|\geq 0.24$. ∎
4.3 Bounds on the Second-Order Partials
Given $S\in\mathcal{B}$, for a triangle $(i,j,k)\in F$, recall our notation $V=X_{j}-X_{i}$, $W=X_{k}-X_{i}$, $z_{i}$ the $z$-coordinate of $X_{i}$, and $\theta_{ijk}$ the angle at $X_{i}$. Throughout this section, we write
$u=\langle V,W\rangle_{X_{i}};\hskip 5.0ptv=\|V\|_{X_{i}};\hskip 5.0ptw=\|W\|_{X_{i}};\hskip 5.0pt\theta=\theta_{ijk}=\arccos(u(vw)^{-1}).$
We remind the readers of the bounds we obtained from $\S$4.2:
  • $\|X_{i}\|\leq 0.8$ ($S\subset B_{0.8}(0)$);
  • $\|V\|,\|W\|\leq 1.6$ ($S\subset B_{0.8}(0)$);
  • $v,w\in[0.5,13]$ (Corollary 4.8);
  • $|\sin\theta|\geq 0.24$ (Corollary 4.12);
  • $|z_{i}|\leq 0.8$ ($S\subset B_{0.8}(0)$).
We remark that the Cauchy-Schwarz inequality further implies $|u|\leq 1.6^{2}$.
Lemma 4.13.
The following bounds on the first-order partials hold for all $l=i,j,k$:
  1. 1.
    $|\partial_{l}u|\leq 10^{3}$;
  2. 2.
    $|\partial_{l}v|,|\partial_{l}w|\leq 10^{3}$;
  3. 3.
    $|\partial_{l}\theta|\leq 10^{7}$.
Proof.
To see $|\partial_{l}u|\leq 10^{3}$, Equation (14) and (15) imply
$|\partial_{i}u|\leq\frac{4\cdot 0.8}{1-0.8^{2}}+\frac{8\cdot 0.8\cdot 1.6^{2}}{0.36^{2}}+\frac{4\cdot 0.8\cdot 1.6^{4}}{0.36^{3}}\leq 10^{3};\hskip 10.0pt|\partial_{l}u|\leq\frac{2\cdot 0.8}{0.36}+\frac{0.8\cdot 1.6^{2}}{0.36^{2}}\leq 10^{3}$
for $l=j$ and $k$. Here, the denominator $1-\langle X_{i},X_{i}\rangle$ is bounded below by $1-0.8^{2}=0.36$.
To see $|\partial_{l}v|\leq 10^{3}$, we first observe that $(2v)^{-1}\leq(2\cdot 0.5)^{-1}=1$. Then, Equation (16) and (17) imply
$|\partial_{i}v|\leq\frac{4\cdot 0.8}{0.36}+\frac{8\cdot 0.8\cdot 1.6^{2}}{0.36^{2}}+\frac{4\cdot 0.8\cdot 1.6^{4}}{0.36^{3}}\leq 10^{3};\hskip 10.0pt|\partial_{j}v|\leq\frac{4\cdot 0.8}{0.36}+\frac{2\cdot 0.8\cdot 1.6^{2}}{0.36^{2}}\leq 10^{3}.$
Moreover, we have $\partial_{k}v=0$. This shows $|\partial_{l}v|\leq 10^{3}$ for all $l$. The argument for $|\partial_{l}w|\leq 10^{3}$ is symmetric, so we will omit it.
Finally, to see $|\partial_{l}\theta|\leq 10^{7}$, Equation (18) implies
$|\partial_{l}\theta|=\frac{\partial_{l}(vw)\cos\theta-\partial_{l}u}{vw\sin\theta}\leq\frac{2\cdot 1.6\cdot 10^{3}+10^{3}}{0.5^{2}\cdot 0.24}\leq 3.2\cdot 10^{6}\leq 10^{7}.$
This completes the proof. ∎
Lemma 4.14.
The following bounds on the second-order partials hold for all $l,m\in\{i,j,k\}$:
  1. 1.
    $|\partial_{lm}u|\leq 10^{4}$;
  2. 2.
    $|\partial_{lm}v|,|\partial_{lm}w|\leq 10^{7}$.
Proof.
To ease notation we write
$a=1-\|X_{i}\|^{2};\hskip 5.0ptt_{1}=\langle X_{i},V\rangle;\hskip 5.0ptt_{2}=\langle X_{i},W\rangle;\hskip 5.0ptt_{3}=\langle V,W\rangle;\hskip 5.0ptt_{4}=\|V\|^{2}.$
From Cauchy-Schwarz inequality, all the $t_{i}$’s satisfy $|t_{i}|\leq 1.6^{2}$, and $a>1-0.8^{2}=0.36$.
We start by computing the second-order partials. Equation (14) and (15) gives us
$\partial_{ii}u=$ $\frac{2}{a}+\frac{2(t_{3}-t_{1}-t_{2}+8z_{i}^{2}-4z_{i}z_{j}-4z_{i}z_{k}+z_{j}z_{k})}{a^{2}}$
$+\frac{4(t_{1}t_{2}+2t_{1}z_{i}z_{k}-4t_{1}z_{i}^{2}+2t_{2}z_{i}z_{j}-4t_{2}z_{i}^{2}+2t_{3}z_{i}^{2})}{a^{3}}+\frac{24t_{1}t_{2}z_{i}^{2}}{a^{4}};$
$\partial_{ij}u=\frac{-1}{a}+\frac{t_{2}-4z_{i}^{2}+3z_{i}z_{k}}{a^{2}}+\frac{4t_{2}z_{i}^{2}}{a^{3}};\hskip 15.0pt\partial_{jk}u=\frac{1}{a}+\frac{z_{i}^{2}}{a^{2}};$
and $\partial_{jj}u=\partial_{kk}u=0$. $\partial_{ik}u$ follows from symmetry.
For the second-order partials of $v$, we have
$\partial_{ii}v=\frac{1}{av}+\frac{t_{4}-2t_{1}+8z_{i}^{2}-8z_{i}z_{j}+z_{j}^{2}}{a^{2}v}+\frac{2(2t_{4}z_{i}^{2}+4t_{1}z_{i}z_{j}-8t_{1}z_{i}^{2}+t_{1}^{2})}{a^{3}v}+\frac{12t_{1}^{2}z_{i}^{2}}{a^{4}v}-\frac{(\partial_{i}v)^{2}}{v};$
$\partial_{ij}v=\frac{-1}{av}+\frac{3z_{i}z_{j}-4z_{i}^{2}+t_{1}}{a^{2}v}+\frac{4t_{1}z_{i}^{2}}{a^{3}v}-\frac{\partial_{i}v\partial_{j}v}{v};\hskip 15.0pt\partial_{jj}v=\frac{-(\partial_{j}v)^{2}}{v}+\frac{1}{av}+\frac{z_{i}^{2}}{a^{2}v};$
and all other partials vanish.
Next, we find bounds on all these partials. The partials $|\partial_{lm}u|$ are bounded above by $|\partial_{ii}u|$, where
$|\partial_{ii}u|\leq\frac{2}{0.36}+\frac{6\cdot 1.6^{2}+34\cdot 0.8^{2}}{0.36^{2}}+\frac{4\cdot 1.6^{4}+56\cdot 1.6^{2}\cdot 0.8^{2}}{0.36^{3}}+\frac{24\cdot 1.6^{4}\cdot 0.8^{2}}{0.36^{4}}\leq 10^{4}.$
The partials $|\partial_{lm}v|$ are bounded above by $\partial_{ii}v$, where
$|\partial_{ii}v|\leq\frac{1}{0.36\cdot 0.5}+\frac{3\cdot 1.6^{2}+17\cdot 0.8^{2}}{0.36^{2}\cdot 0.5}+\frac{24\cdot 1.6^{2}\cdot 0.8^{2}+2\cdot 1.6^{4}}{0.36^{3}\cdot 0.5}+\frac{12\cdot 1.6^{4}\cdot 0.8^{2}}{0.36^{4}\cdot 0.5}+\frac{10^{6}}{0.5}\leq 10^{7}.$
The argument for $|\partial_{lm}w|\leq 10^{7}$ is symmetric, so we will omit it. ∎
Proof of Lemma 4.4.
From Equation (18), we compute $\partial_{lm}\theta$ as
$\partial_{lm}\theta$ $=\frac{\partial_{lm}(vw)\cos\theta-\partial_{l}(vw)\sin\theta\,\partial_{m}\theta-\partial_{lm}u}{(v^{2}w^{2}-u^{2})^{1/2}}-\frac{(\partial_{l}(vw)\cos\theta-\partial_{l}u)\,\partial_{l}(v^{2}w^{2}-u^{2})}{2(v^{2}w^{2}-u^{2})^{3/2}}$
$=\frac{\partial_{lm}(vw)\cos\theta-\partial_{l}(vw)\sin\theta\,\partial_{m}\theta-\partial_{lm}u}{vw\sin\theta}-\frac{(\partial_{l}(vw)\cos\theta-\partial_{l}u)\,\partial_{l}(v^{2}w^{2}-u^{2})}{2v^{3}w^{3}\sin^{3}\theta}.$
Then, from $|\sin\theta\geq 0.24|$, $u\leq 1.6^{2}$, and $v,w\in[0.5,13]$ we know that $|\partial_{lm}\theta|$ is bounded above by
$\frac{2\cdot 10^{6}+2\cdot 13\cdot 10^{7}+2\cdot 13\cdot 10^{10}+10^{4}}{0.5^{2}\cdot 0.24}+\frac{(2\cdot 13\cdot 10^{3}+10^{3})\cdot(4\cdot 13\cdot 10^{3}+2\cdot 13^{2}\cdot 10^{3})}{2\cdot 0.5^{6}\cdot 0.24^{3}},$
which is less than $10^{14}$. This completes the proof. ∎